HUBBLE SPACE TELESCOPE DAILY REPORT #2286 PERIOD COVERED: 0000Z (UTC) 01/13/99 - 0000Z (UTC) 01/14/99 Daily Status Report as of 014/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed STIS/CCD 7672 (STIS CCD Sensitivity Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star AGK+81D266 in order to monitor the sensitivity of each CCD grating mode to detect any change due to contamination or other causes. A wavelength calibration image was also taken. The observations were completed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. As described in yesterday's report and HSTAR 6897, the acquisition for the first iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting two observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.4 Completed WF/PC-2 6554 (Extragalactic Globular Cluster Systems) The WF/PC-2 was used to make observations of a extragalactic globular cluster system which is located in galaxy NGC-1023. This is part of a survey which will cover a range in galaxy luminosity, morphological type, environment, galactic cluster specific frequency and galactic cluster mean metallicity. As described in yesterday's report and HSTAR 6897, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting two observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.5 Completed Two Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.6 Completed STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was to be used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. As described in HSTAR 6901 and 2.1, the acquisition for this proposal failed and the take data flag remained down. 1.7 Completed WF/PC-2 7376 (Understanding Halo Microlensing and Variable Stars in M31: 1} Snapshot Survey of Candidates) The WF/PC-2 was to be used to make observations of the galaxy M-31-Lens system #3 to help understand its halo microlensing and Variable Stars. Six microlensing candidates affecting stars in M31 and presumably due to sub-solar-mass MACHOs are being studied. Simultaneously we can explore the stellar population in the relatively unexplored disk/bulge transition via its variable stars, luminosity function and color-magnitude diagram. As described in HSTAR 6901 and 2.1, the acquisition for this proposal failed and the take data flag remained down. Two observations were affected. 1.8 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was to be used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. As described in HSTAR 6901 and 2.1, the acquisition for this proposal failed and the take data flag remained down. 1.9 Completed STIS/CCD/MA1 7737 (Low Redshift Lyman Alpha Forest Clouds: Their Kinematics and Their Relation to Galaxies) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make spectrographic observations of background quasar QSO-123050 in order to obtain foreground low redshifted Lyman-Alpha observations of the forest clouds. We should be able to shed light on the question of whether the low redshift Lyman-Alpha forest is related to galaxies or represents a truly intergalactic phenomenon. A wave length calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.10 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II parallel archive proposal continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 5 As documented in HSTAR 6901, the acquisition scheduled for 013/141433Z failed when the FGS-3 scan steps limits was exceeded. The proposals described in 1.6, 1.7, and 1.8 were affected. Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 11 Successful : 11 The roll delay update scheduled for 013/1046Z failed completely. HSTAR 6898 was written. The subsequent acquisition, however, was successful. 2.3 Operations Notes: The SSR EDAC error counter was cleared three times per ROP SR-1. A NSSC-1 status buffer dump was obtained via CCS at 013/1835Z. ROP NS-3 was used. A period of CCS comanding was completed nominally at 013/1907Z. Using ROP IC-2, the SSA transmitter was turned on and off around an SSR engineering playback. FHST performance data was gathered at 014/0210. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH